A 5.00-L air sample has a pressure of 107 kPa at a temperature of –50.0ºC. If the temperature is raised to 102ºC and the pressure drops to 60.0 kPa, what will the new volume be?


1
Expert's answer
2021-02-25T03:41:31-0500

There are two stages of gas:


At first stage

"V_1=" 5 L

"P_1""=" 107kPa

"T_1=" "-50" degre cecsius= "-" 50"+" 273"=" 223 K


At second stage:

"P_2" =60kPa

"T_2=" 102 degre celsius = 102"+" 273"=" 375 K

"V_2=V_2" (let)


From gas eqn "\Rightarrow" PV "=" nRT

"\Rightarrow" n "=\dfrac{PV}{RT}"


So here in question moles are constant at both stages,so:


"\Rightarrow" "n_1=n_2"

"\Rightarrow" "\dfrac{P_1V_1}{RT_1}=\dfrac{P_2V_2}{RT_2}"

R is conatant from both the sides so


"\Rightarrow" "V_2=\dfrac{P_1V_1T_2}{T_1P_2}"


"\Rightarrow" "V_2" "=\dfrac{107\times5\times375}{223\times60}" "=" 14.99 L"=" 15.00 L approximtely